# Melody Is All You Need For Music Generation

### v3

@@ 6.3. Satisfaction
We conducted a survey with 125 randomly selected participants who do not have specialized knowledge in music. This questionnaire, tailored for general users, included questions about the clarity, infectivity, and freshness of the generated music. More details on the questionnaire can be found in Appendix A.2. The results, presented in Table 4, show that, on average, 17.71% and 42.94% of respondents rated the generated music as ’Very Consistent’ and ’Consistent,’ respectively, with the desired attributes of clarity, infectivity, and freshness. The overall score of 3.54 further reflects that most participants were generally satisfied with the music generated by $\text{MG}^{2}$.
@@ 6.4. Quality
To assess the quality of the generated music, we gathered responses from 18 musicians with formal music education. These musicians answered professional questions tailored to evaluate the music’s technical and artistic quality. Detailed descriptions of these questions can be found in Appendix A.3. As shown in Table 5, more than half of the musicians, represented by averages of 12.80 and 42.76 in the final row, rated the generated music as high quality. This feedback underscores the professional viability of $\text{MG}^{2}$’s outputs.
@@ 6.5. Market Potential
We evaluate the market potential of $\text{MG}^{2}$ in real creation scenarios by surveying 20 short video bloggers from popular social video platforms in China, including TikTok, REDnote, and Bilibili. These bloggers have fan counts ranging from 1,000 to over 20,000. The questionnaire, detailed in Appendix A.4, aimed to assess their willingness to use the generated music. The music used in earlier sections was evaluated on its feasibility for short video editing and the bloggers’ willingness to pay for it. The results, presented in Tables 6 and 7, show that 75.72% of bloggers (combining Very Consistent and Consistent responses) believe the music is feasible for short video editing. Additionally, 46.43% of bloggers expressed a willingness to pay for the generated music. We also evaluated the usefulness from another perspective by asking bloggers whether they would use the generated music in their content creation. We presented three types of music—exciting, comical, and suspenseful music—each with three clips. Bloggers were asked to select zero to three clips they would use in their creations. As shown in Table 8, over 90% of bloggers indicated they would use at least one music clip for their content. These results, reflected in the three tables, demonstrate the practical value of $\text{MG}^{2}$ from multiple perspectives, highlighting its potential in real-world applications such as video creation.
@@ 8. Acknowledgements
Appendix A Human Evaluation Measure
@@ A.3. Musican
A.4. Short Video Blogger

### v4

@@ 6.3. Satisfaction
We conducted a survey with 124 randomly selected participants who do not have specialized knowledge in music. This questionnaire, tailored for general users, included questions about the clarity, infectivity, and freshness of the generated music. More details on the questionnaire can be found in Appendix B.2. The results, presented in Table 4, show that, on average, 17.71% and 42.94% of respondents rated the generated music as ’Very Consistent’ and ’Consistent,’ respectively, with the desired attributes of clarity, infectivity, and freshness. The overall score of 3.54 further reflects that most participants were generally satisfied with the music generated by $\text{MG}^{2}$.
@@ 6.4. Quality
To assess the quality of the generated music, we gathered responses from 18 musicians with formal music education. These musicians answered professional questions tailored to evaluate the music’s technical and artistic quality. Detailed descriptions of these questions can be found in Appendix B.3. As shown in Table 5, more than half of the musicians, represented by averages of 12.80 and 42.76 in the final row, rated the generated music as high quality. This feedback underscores the professional viability of $\text{MG}^{2}$’s outputs.
@@ 6.5. Market Potential
We evaluate the market potential of $\text{MG}^{2}$ in real creation scenarios by surveying 20 short video bloggers from popular social video platforms in China, including TikTok, REDnote, and Bilibili. These bloggers have fan counts ranging from 1,000 to over 20,000. The questionnaire, detailed in Appendix B.4, aimed to assess their willingness to use the generated music. The music used in earlier sections was evaluated on its feasibility for short video editing and the bloggers’ willingness to pay for it. The results, presented in Tables 6 and 7, show that 75.72% of bloggers (combining Very Consistent and Consistent responses) believe the music is feasible for short video editing. Additionally, 46.43% of bloggers expressed a willingness to pay for the generated music. We also evaluated the usefulness from another perspective by asking bloggers whether they would use the generated music in their content creation. We presented three types of music—exciting, comical, and suspenseful music—each with three clips. Bloggers were asked to select zero to three clips they would use in their creations. As shown in Table 8, over 90% of bloggers indicated they would use at least one music clip for their content. These results, reflected in the three tables, demonstrate the practical value of $\text{MG}^{2}$ from multiple perspectives, highlighting its potential in real-world applications such as video creation.
@@ 8. Acknowledgements
Appendix A Accurate Alignment
We present the text descriptions of the used pairs in Figure 2 as follows:
  • (a) This audio contains someone playing a piece on cello ranging from the low register up into the higher register. This song may be playing during a live performance.
  • (b) The song is an instrumental. The tempo is medium with a guitar playing a romantic lead, steady drumming, rock drumming, percussive bass line, cymbals crashing and guitar strumming rhythm. The song is exciting and youthful.
  • (c) This instrumental song features a violin playing the main melody. This is accompanied by a guitar playing chords. The piano plays backing chords. At the end of the song, the piano plays a fill. The bass is played on a double bass. There is no percussion in this song. The mood of this song is romantic. This song can be played in a romantic movie.
  • (d) This music is instrumental. The tempo is fast with a spirited cello and harpsichord accompaniment. The music is lively, intense, rich, complex , exquisite and engaging. This music a Western Classical.
Appendix B Human Evaluation Measure
@@ B.3. Musican
B.4. Short Video Blogger